Circular No. 6432

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRS 1915+105 E. Gérard, Observatoire de Paris-Meudon, reports that a new radio

outburst was detected from this hard x-ray transient in Aquila with the Nançay Radio Telescope on July 9. The observed radio fluxes were: July 8.009 UT, < 26 mJy at 1414 MHz and < 40 mJy at 3310 MHz; July 9.009, < 30 mJy at 1414 MHz and 83 mJy at 3310 MHz; July 9.995, 140 mJy at 1414 MHz and 94 mJy at 3310 MHz. Multi-wavelength observations are encouraged.

2S 1417–624 M. H. Finger and M. Scott, Universities Space Research Association;

T. A. Prince and B. Vaughan, California Institute of Technology; and R. B. Wilson, Marshall Space Flight Center, NASA, report for the Compton Observatory BATSE team: “Renewed activity from the x-ray-transient pulsar 2S 1417–624 is being observed. Hard x-ray pulses were first detected on June 30, with the pulsed flux rising to 17 ± 3 mCrab (20–50 keV) by July 8. A barycentric frequency of 0.05694816(25) Hz on July 4.5 UT and barycentric frequency rate of 3.35(14) × 10^{−11} Hz/s were obtained from a fit of frequency measurements from July 1–8. The source had been undetected by BATSE since a series of six outbursts were observed, beginning on 1994 Aug. 26 (IAUC 6075) and ending 1995 July 8.”

CAL 83 P. Kahabka, University of Amsterdam, reports: “During ROSAT/HRI

observation on Apr. 28.162–28.222 UT, the supersoft source numbered 83 by Long et al. (1981, ApJ 248, 925; = 1E 0543.8–6823) was not detected. On Apr. 7, the ROSAT/HRI detected this source at a rate of 0.2 count/s. The x-ray off-state on Apr. 28 implies a reduction in x-ray flux by a factor of > 43 in 21 days. High- and low-intensity states of CAL 83 are also seen in the ultraviolet. CAL 83 may be similar to the recurrent LMC transient RX J0513.9–6951, which shows optical dips with a recurrence period of ∼ 100–200 days during which x-ray outbursts occur. The disappearance of the source in x-rays is understood to be due to a shift of the temperature to lower values, rendering the source undetectable in x-rays. The length of x-ray on- and off-states may be different in RX J0513.9–6951 and CAL 83, due to different white-dwarf masses and mass-accretion rates.”

1996 July 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green